6. Find the particular part of the solution of the difference equation y(n+2) – 2y(n+1)+y(n) = 4 for n <0.
Answer #1

y (n+2) - 2 4h+1) + yn)=4 for no yh [y? - 2y + 1]y = 4 Put nzo [4² - 2y + 1] 4 = 4 il y = D ie du [2² - 2D + 1] y=4 92 02-20+